OpenStudy (anonymous):

Solve for t. 792πt=2π

zepdrix (zepdrix):

Hmmm so we want to solve for t. t is being multiplied by a bunch of stuff. We can isolate the t by dividing by the stuff.\[\Large 792\pi t\quad=\quad 2\pi\quad\to\qquad t\quad=\quad \frac{2\pi}{792\pi}\]Then we just simplify! :)

OpenStudy (anonymous):

so that would be 1/396?